OpenStudy (anonymous):

Evaluate the polynomial 6x - 7 for... 1/2n2 -1/2n and y = 2

OpenStudy (anonymous):

HOW to EVALUATE POLYNOMIAL Evaluate a^2b for a = –2, b = 3, c = –4, and d = 4. To find my answer, I just plug in the given values, being careful to use parentheses, particularly around the minus signs: Copyright © Elizabeth Stapel 2000-2011 All Rights Reserved \[\huge (–2)^2 (3) = (4)(3) = 12 \]

OpenStudy (anonymous):

is your doubt clear @kmstevens10

OpenStudy (anonymous):

sorry after that it asks How many segments can you draw if you mark 8 points on the circle?

OpenStudy (anonymous):

@Muskan as u r here are u educated bout this topic if not i will continue

OpenStudy (anonymous):

wait nevermind forget that last part! that was a seperate question im confusing myself :)

OpenStudy (anonymous):

A. 15 B. –20 C. 4 D. –16 these are my choices

OpenStudy (anonymous):

hmmm k,....

OpenStudy (anonymous):

its 4

OpenStudy (anonymous):

thanku

OpenStudy (anonymous):

:))